OpenStudy (anonymous):

Simplify sin^2(x-1)/cos(-x)

OpenStudy (anonymous):

@Hero

hero (hero):

I know what happened now

hero (hero):

You posted: \[\frac{\sin^2(x - 1)}{-\cos(x)}\] But what you really meant to post was \[\frac{\sin^2(x) - 1}{\cos(-x)}\]

hero (hero):

Those are two different expressions

OpenStudy (anonymous):

ah! did i type it in wrong?!

hero (hero):

The second one simplifies to one of your answer choices.

OpenStudy (anonymous):

i'm sorry!

hero (hero):

Yes, you did

hero (hero):

I was a bit thrown off at first, but I assumed what you wrote was correct.

hero (hero):

Well, in this case \(\sin^2x = 1 - \cos^2(x)\)

hero (hero):

That should be the only hint you need.

OpenStudy (anonymous):

\[\frac{ 1-\cos ^2\left( x-1 \right) }{ \cos \left( -x \right) }\]

hero (hero):

\[\frac{ 1-\cos ^2( x)-1 }{ \cos \left( -x \right) }\]

hero (hero):

Why do you keep putting the 1 in parentheses?

hero (hero):

Is that what you see on your reference sheet?

OpenStudy (anonymous):

but isn't it apart of it?

hero (hero):

No, it isn't.

OpenStudy (anonymous):

it looks like this

hero (hero):

Write it out on the board what you see

hero (hero):

If it has the 1 in parentheses, then it is a mistake.

OpenStudy (anonymous):

\[\frac{ \sin ^2x-1 }{ \cos \left( -x \right) }\]

OpenStudy (anonymous):

ahhh the sin^2x is seperate from the 1

OpenStudy (anonymous):

ahhh

hero (hero):

Yes, and you only replace the \(\sin^2(x)\) with \(1 - \cos^2(x)\)

OpenStudy (anonymous):

1 - (-1) = 2, so would it be 2 .... cos^2(x)

OpenStudy (anonymous):

so would it be -cos^2(x)/cos(-x)

hero (hero):

Hmmm, but hang on.

OpenStudy (anonymous):

because the problem became:

hero (hero):

We know what the end result should be

hero (hero):

1 - 1 = 0 like I said

OpenStudy (anonymous):

\[\frac{ 1-\cos ^2\left( x \right)-1 }{ \cos \left( -x \right) }\]

hero (hero):

Yes, that is correct now.

OpenStudy (anonymous):

haha i'm sorry

OpenStudy (anonymous):

so it the top cos going to be negative stil?

hero (hero):

And then \[\dfrac{-\cos^2(x)}{\cos(-x)}\] is what you should have up to this point.

hero (hero):

But \(\cos(-x) = -\cos(x)\)

hero (hero):

However there's still an issue somewhere.

hero (hero):

Because \[\dfrac{-\cos^2x}{\cos(-x)} = \dfrac{-\cos^2(x)}{-\cos(x)}\]

OpenStudy (anonymous):

\[\frac{ -\cos ^2\left( x \right) }{ -\cos \left( x \right) }\]

hero (hero):

If we simplify that we end up with only \(\cos(x)\) instead of \(-\cos(x)\)

OpenStudy (anonymous):

one of the choices are cos(x)

hero (hero):

I know, however, my calc is never wrong.

OpenStudy (anonymous):

haha okay, so it shouldn't be cos(x) but -cos(x)

hero (hero):

Hang on. Let me check something

OpenStudy (anonymous):

hmm i'm trying to think of how to help but my calc is usually wrong lol

hero (hero):

I have it now. In the denominator, \(\cos(-x) = \cos(x)\). So there we go.

hero (hero):

So to be clear, the correct answer is \(-\cos(x)\)

OpenStudy (anonymous):

yes, that makes sense! thank you, i think i am starting to kind of understand but its very hard stuff. I do have another one, could i open another question and you can help me with it?

hero (hero):

Okay, maybe this next one will be better.

OpenStudy (anonymous):

haha you were perfect, thank you very much!